OpenStudy (anonymous):

use a system of equations in two variables to solve the following problem: The sum of two numbers is 64. One number is 14 more than the other. Find the numbers?

OpenStudy (anonymous):

x + y = 64 x = y + 14

OpenStudy (anonymous):

I think I got it from here

OpenStudy (anonymous):

y +14= 64 y= 50

OpenStudy (anonymous):

I should get 39, 25